# Universal Transient Stability Analysis: A Large Language Model-Enabled Dynamics Prediction Framework

### v1

@@ 3.3.3 Schedule Sampling Phase
To mitigate the previously described training-inference mismatch, a scheduled sampling fine-tuning stage is proposed. Following a curriculum learning strategy [lamb2016professor], SchS focuses on the top-$K$ error-prone trajectories, denoted by $x_{i}(t)\in D_{\text{hard}}$ for $t\in[1,T]$ identified within $D_{\text{train}}$, as identified in Algorithm 1. Critically, by emulating the iterative, self-conditioned prediction of online deployment, this process compels the model to develop robustness against its own error accumulation, thereby improving long-horizon iterative prediction in online TSA tasks.
During SchS stage, given an initial input sequence $x_{i,\text{input}}\in\mathbb{R}^{L_{\text{seq}}}$ spanning time steps $1:L_{\text{seq}}]$ from trajectory $x_{i}$, the model first generates a prediction for the next time window $L_{\text{seq}}+1,L_{\text{seq}}+L_{\text{pred}}]$:
$\hat{\mathbf{x}}_{i,\text{target}}=\mathbf{F}\left(x_{i,\text{input}}\right)\in\mathbb{R}^{L_{\text{pred}}}.$ (19)
To generate the subsequent prediction, a new input sequence, $x_{i,\text{input}}^{\text{next}}$, is constructed by sliding the window forward. The core of the SchS strategy lies in how this new input is assembled. It consists of a known segment from the original ground-truth trajectory ($x_{i,\text{orig}}$) and a ”mixed” segment ($x_{i,\text{mixed}}$), which is probabilistically sampled from either the ground-truth target ($x_{i,\text{target}}$) or the model’s own previous prediction ($\hat{x}_{i,\text{target}}$):
@@
where $E_{\text{start}}$ is the epoch at which decay begins, and $E_{\text{max}}$ is the maximum training epoch in SchS stage. The fine-tuning loss, $\mathcal{L}_{\mathrm{SchS}}$, aggregates the MSE across the predicted trajectory sequence $\hat{x}_{i}(t)$:
@@
8:   for $j=1,2,\ldots,\lfloor(T-L_{\text{seq}})/L_{\text{pred}}\rfloor$ do
@@ 3.4 Online Application
While channel independence processes each of the $n_{x}$ state variables separately in the offline fine-tuning stage, we mitigate potential computational overhead in online applications by treating the $n_{x}$ channels as a unified mini-batch to exploit GPU parallelization and concurrent predict of all state variables. The online prediction begins with an initial multivariate observation from PMUs, denoted as $\textbf{x}^{(0)}_{\text{input}}\in\mathbb{R}^{n_{\textbf{x}}\times L_{\text{seq}}}$. This multi-channel input is decomposed into $n_{x}$ independent univariate patch sequences $\mathbf{x}_{\text{patch}}^{(0)}\in\mathbb{R}^{P\times L_{P}}$ as described in subsection 3.1, which are then stacked along the batch dimension via $\operatorname{Batch}(\cdot)$, forming $\textbf{x}^{(0)}_{\text{batch}}\in\mathbb{R}^{n_{\textbf{x}}\times P\times L_{p}}$. The TSA-LLM then predicts the target sequence of all state variables $\hat{\textbf{x}}^{(0)}_{\text{target}}\in\mathbb{R}^{n_{\textbf{x}}\times L_{\text{pred}}\times 1}$. Then, iterative prediction for each step $j$ is performed sequentially as follows:
@@
where $\oplus$ denotes the sliding window update mechanism, which discards the oldest input segment and appends the latest prediction, and $\operatorname{Preprocess}(\cdot)$ follows the data processing pipeline in subsection 3.1. This fully iterative online prediction aligns with the SchS strategy where the sampling probability $\epsilon_{k}$fixed at 0, forcing the model to rely entirely on its own generated trajectory.
@@ 3.5 Performance Indices Evaluation
where ${D}_{\mathrm{C}}$ is the test set for category C, $|{D}_{\mathrm{C}}|$ is the number of trajectories within it, and $T_{\text{pred}}=T-L_{\text{seq}}$ is the prediction length. The terms $\hat{x}_{i}(t)$ and $x_{i}(t)$ denote the predicted and ground-truth values, respectively, for the $i$-th dynamic trajectories at time $t$. This categorized analysis allows us to not only quantify but also explain the performance disparities observed in its handling of stable versus unstable system dynamics.
@@ 4 Case study
The proposed TSA-LLM framework is rigorously evaluated on the New England 39-bus [zhao2022structure] and modified Iceland 189-bus [cai2020data] systems. On the 39-bus system, we assess the framework’s zero-shot generalization to mixed-stability OCs and unseen faults and conduct a representation analysis of its unstable OC handling mechanism. The 189-bus system is applied to validate the model’s few-shot scalability on a more complex, heterogeneous grid. The evaluation is completed with analyses of training/inference costs, an ablation study of model components, and a final qualitative demonstration of the predictive enhancements gained from using frozen T-blocks.
@@ 4.1.1 Test System and Simulation Settings
Two standard test systems were applied: the New England 39-bus system (10 SGs, 34 lines) and a modified Iceland 189-bus system (33 SGs, 206 lines). The latter was adapted for 20% RES penetration by replacing two SGs with wind units whose speeds were sampled from a Weibull distribution [lu2024advanced]. Simulations includes N1N-1, N2N-2 and N3N-3 faults for the 39-bus system and N1N-1 faults for the 189-bus system. For each contingency, simulations are conducted as follows: 1) All system loads were randomly varied between 70% and 130% of their nominal values; 2) An optimal power flow was solved using MATPOWER to establish the pre-fault OCs; 3) All SGs were represented by the fourth-order model; 4) Each contingency was simulated by applying the specified number of concurrent three-phase-to-ground faults to random lines at t=1.0st=1.0~s and cleared after a random duration of up to 0.3 s; 5) The post-fault dynamic response was simulated for 10 seconds with a 0.02 s time step by Power System Analysis Toolbox.
@@ 4.1.2 Dataset Organization
For fine-tuning and validation, a primary dataset of 9,000 trajectories (rotor angles and speeds) was generated from N1N-1 contingencies on the 39-bus system, covering a mix of stable and unstable trajectories. This was partitioned into 8,000 training ($D_{\text{train}}$) and 1,000 validation ($D_{\text{val}}$) trajectories. Each trajectory was then segmented via a sliding window ($L_{\text{seq}}=65,L_{\text{pred}}=1$) to yield approximately $1.8\times 10^{7}$ training and $2\times 10^{6}$ validation samples. Moreover, a 3,300-trajectory set ($D_{\text{train}}^{189\text{bus}}$) generated from the 189-bus system was used for cross-system few-shot fine-tuning. Model performance was evaluated against four distinct test sets, each designed to assess a specific aspect of universality. The details of these datasets are summarized in Table 2. The subsequent analysis primarily presents rotor angle predictions and results for other variables like rotor speeds are analogous and detailed in subsection 4.6.
@@ 4.1.3 Model Implementation Details
The proposed TSA-LLM, based on the GPT architecture (124M parameters, 12 layers, 12 heads, 768-dim embedding), was fine-tuned for up to 10 epochs using a two-stage TeaF and SchS scheme. Training utilized the Adam optimizer with a $1\times 10^{-4}$ initial learning rate managed by a cosine annealing scheduler and an early stopping criterion. To evaluate its performance, TSA-LLM was benchmarked against three baseline models. The first two are the DNR framework from [zhao2022structure] and a standard LSTM network as presented in [ye2024use], for which we employed the hyperparameter settings as reported in their respective publications. The third baseline is an Encoder-only Transformer (ENC), constructed based on [vaswani2017attention] to highlight performance differences arising from different attention mechanisms. This ENC model employs the bidirectional attention mechanism and is configured with 3 transformer layers, 6 bidirectional attention heads, and a 128-dim embedding. All baselines were implemented for multivariate dynamics prediction in PyTorch and trained on NVIDIA A100 GPU.
@@ 4.2.1 Performance Comparison
This section evaluates the model’s robustness against mixed stable or unstable OCs on the $D_{\text{test}}^{N-1}$, which is collected in $\{\text{S},\text{U},\text{H}\}$ situations as discussed in subsection 3.5. The quantitative results are summarized in Table 3, where bold and underlined values denote the best and second-best performance, respectively. As shown, TSA-LLM substantially outperforms all baseline models across nearly all metrics, with the performance gap being most pronounced in challenging unstable and hybrid cases. While traditional methods like DNR and LSTM struggle significantly in these scenarios with DNR’s $MSE_{U}$ value even exceeding 100, TSA-LLM achieves remarkable error reductions. Specifically, TSA-LLM reduces $MAE_{U}$, $MSE_{U}$, $MAE_{H}$, and $MSE_{H}$ by at least 84.49%, 97.27%, 83.49%, and 97.68%, respectively, compared to baselines. These quantitative findings are further corroborated by visualization of the rotor angle trajectories predicted by TSA-LLM in Figure 3, which show a close tracking of the ground truth. This significant disparity in performance can be attributed to fundamental architectural differences. Traditional RNN-based models (LSTM and DNR) suffer from vanishing gradient issue and exhibit constrained model capacity, which fundamentally limits their ability to capture complex nonlinear dynamics in unstable trajectory prediction. In contrast, the transformer architecture, utilized by both our proposed TSA-LLM and the ENC baseline, is far more effective at capturing long-range temporal dependencies, resulting in demonstrably superior accuracy in these challenging scenarios.
4.2.2 Explaining Robust Prediction in Unstable Scenarios
To further investigate the source of TSA-LLM’s robustness, we conducted an interpretive analysis focused on its feature representations of OOS dynamics, which are critical in unstable scenarios. We selected a typical unstable case from the 39-bus system where SG $\delta_{5}$ loses synchronism (Figure 4(a)). Final-layer feature representations for the OOS SG ($\delta_{5}$) and two stable SGs ($\delta_{1},\delta_{10}$) were then extracted and projected into 2D space from both TSA-LLM and an identically trained channel-independent LSTM model for visualization with t-SNE [zhou2023one]. The t-SNE visualizations reveal a stark contrast between the two architectures. For TSA-LLM (Figure 4(b)), features from $\delta_{5}$ form a distinct and clearly separable cluster, indicating effective learning of instability characteristics. However, the LSTM model exhibits substantial mixing of these representations (Figure 4(c)). This demonstrates the inherent difficulty of RNN architectures in distinguishing the unique behavior of OOS generators, which correlates with their lower prediction accuracy and underscores the superior feature learning of the proposed attention-based architecture.
@@ 4.2.2 Explaining Robust Prediction in Unstable Scenarios
(b) Actual $\delta$ in unstble OC
(c) Predicted $\delta$ in stable OC
(d) Predicted $\delta$ in unstble OC
Figure 3: New England 39-bus system: The evaluation of rotor angle under the stable and unstable OC. The predicted $\delta$ by TSA-LLM ((c) and (d)) demonstrates close alignment with ground truth ((a) and (b)).
Figure 4: The t-SNE visualization of stable/unstable sample feature maps for the proposed TSA-LLM and LSTM.

### v2

@@ 3.3.3 Schedule Sampling Phase
To mitigate the previously described training-inference mismatch, a scheduled sampling fine-tuning stage is proposed. Following a curriculum learning strategy [16], SchS focuses on the top-$K$ error-prone trajectories, denoted by $x_{i}(t)\in D_{\text{hard}}$ for $t\in[1,T]$ identified within $D_{\text{train}}$, as identified in Algorithm 1. This top-$K$ selection, rather than using the full set of trajectories, offers three key advantages: (i) it avoids gradient dilution, since the full set contains many “easy” trajectories where the model already performs well after TeaF, whose gradients would dominate and dilute the signal from hard cases that need improvement [41]; (ii) it blackuces computational cost, as SchS requires full iterative pblackiction over each trajectory and training on the full set would incur substantially higher per-epoch cost; and (iii) it allocates the limited SchS training budget to the trajectories that most impact online deployment, where cumulative error and training-inference mismatch are most severe. Critically, by emulating the iterative, self-conditioned pblackiction of online deployment, this process compels the model to develop robustness against its own error accumulation, thereby improving long-horizon iterative pblackiction in online TSA tasks.
During SchS stage, given an initial input sequence $x_{i,\text{input}}\in\mathbb{R}^{L_{\text{seq}}}$ spanning time steps $1:L_{\text{seq}}]$ from trajectory $x_{i}$, the model first generates a pblackiction for the next time window $L_{\text{seq}}+1,L_{\text{seq}}+L_{\text{pblack}}]$:
$\hat{\mathbf{x}}_{i,\text{target}}=\mathbf{F}\left(x_{i,\text{input}}\right)\in\mathbb{R}^{L_{\text{pblack}}}.$ (19)
To generate the subsequent pblackiction, a new input sequence, $x_{i,\text{input}}^{\text{next}}$, is constructed by sliding the window forward. The core of the SchS strategy lies in how this new input is assembled. It consists of a known segment from the original ground-truth trajectory ($x_{i,\text{orig}}$) and a ”mixed” segment ($x_{i,\text{mixed}}$), which is probabilistically sampled from either the ground-truth target ($x_{i,\text{target}}$) or the model’s own previous pblackiction ($\hat{x}_{i,\text{target}}$):
@@
where $E_{\text{start}}$ is the epoch at which decay begins, and $E_{\text{max}}$ is the maximum training epoch in SchS stage. The fine-tuning loss, $\mathcal{L}_{\mathrm{SchS}}$, aggregates the MSE across the pblackicted trajectory sequence $\hat{x}_{i}(t)$:
@@
8:   for $j=1,2,\ldots,\lfloor(T-L_{\text{seq}})/L_{\text{pblack}}\rfloor$ do
@@ 3.4 Online Application
While channel independence processes each of the $n_{x}$ state variables separately in the offline fine-tuning stage, we mitigate potential computational overhead in online applications by treating the $n_{x}$ channels as a unified mini-batch to exploit GPU parallelization and concurrent pblackict of all state variables. The online pblackiction begins with an initial multivariate observation from PMUs, denoted as $\textbf{x}^{(0)}_{\text{input}}\in\mathbb{R}^{n_{\textbf{x}}\times L_{\text{seq}}}$. This multi-channel input is decomposed into $n_{x}$ independent univariate patch sequences $\mathbf{x}_{\text{patch}}^{(0)}\in\mathbb{R}^{P\times L_{P}}$ as described in subsection 3.1, which are then stacked along the batch dimension via $\operatorname{Batch}(\cdot)$, forming $\textbf{x}^{(0)}_{\text{batch}}\in\mathbb{R}^{n_{\textbf{x}}\times P\times L_{p}}$. The TSA-LLM then pblackicts the target sequence of all state variables $\hat{\textbf{x}}^{(0)}_{\text{target}}\in\mathbb{R}^{n_{\textbf{x}}\times L_{\text{pblack}}\times 1}$. Then, iterative pblackiction for each step $j$ is performed sequentially as follows:
@@
where $\oplus$ denotes the sliding window update mechanism, which discards the oldest input segment and appends the latest pblackiction, and $\operatorname{Preprocess}(\cdot)$ follows the data processing pipeline in subsection 3.1. This fully iterative online pblackiction aligns with the SchS strategy where the sampling probability $\epsilon_{k}$fixed at 0, forcing the model to rely entirely on its own generated trajectory.
@@ 3.5 Performance Indices Evaluation
where ${D}_{\mathrm{C}}$ is the test set for category C, $|{D}_{\mathrm{C}}|$ is the number of trajectories within it, and $T_{\text{pblack}}=T-L_{\text{seq}}$ is the pblackiction length. The terms $\hat{x}_{i}(t)$ and $x_{i}(t)$ denote the pblackicted and ground-truth values, respectively, for the $i$-th dynamic trajectories at time $t$. This categorized analysis allows us to not only quantify but also explain the performance disparities observed in its handling of stable versus unstable system dynamics.
@@ 4 Case study
The proposed TSA-LLM framework is rigorously evaluated on the New England 39-bus system [43] and the modified Iceland 189-bus system [4]. On the New England 39-bus system, we assess the framework’s zero-shot generalization to mixed-stability OCs and unseen faults and conduct a representation analysis of its unstable OC handling mechanism. The Iceland 189-bus system is used to validate the model’s few-shot scalability on a more complex, heterogeneous grid. The evaluation is completed with analyses of training/inference costs, an ablation study of model components, and a final qualitative demonstration of the pblackictive enhancements gained from using frozen T-blocks.
@@ 4.1.1 Test System and Simulation Settings
Two standard test systems were applied: the New England 39-bus system (10 SGs, 34 lines) and the modified Iceland 189-bus system (33 SGs, 206 lines). The latter was adapted for 20% RES penetration by replacing two SGs with wind units whose speeds were sampled from a Weibull distribution [20, 31]. Simulations include N1N-1, N2N-2, and N3N-3 faults for the New England 39-bus system and N1N-1 faults for the Iceland 189-bus system. For each contingency, simulations are conducted as follows: 1) All system loads were randomly varied between 70% and 130% of their nominal values; 2) An optimal power flow was solved using MATPOWER to establish the pre-fault OCs; 3) All SGs were represented by the fourth-order model; 4) Each contingency was simulated by applying the specified number of concurrent three-phase-to-ground faults to random lines at t=1.0st=1.0~s and cleablack after a random duration of up to 0.3 s; 5) The post-fault dynamic response was simulated for 10 seconds with a 0.02 s time step by Power System Analysis Toolbox.
@@ 4.1.2 Dataset Organization
For fine-tuning and validation, a primary dataset of 9,000 trajectories (rotor angles and speeds) was generated from N1N-1 contingencies on the 39-bus system, covering a mix of stable and unstable trajectories. This was partitioned into 8,000 training ($D_{\text{train}}$) and 1,000 validation ($D_{\text{val}}$) trajectories. Each trajectory was then segmented via a sliding window ($L_{\text{seq}}=65,L_{\text{pblack}}=1$) to yield approximately $1.8\times 10^{7}$ training and $2\times 10^{6}$ validation samples. Additionally, trajectories from the 189-bus system were generated to evaluate cross-system few-shot adaptation on a more complex, heterogeneous grid. Model performance was evaluated against four distinct test sets, each designed to assess a specific aspect of universality. The details of these datasets are summarized in Table 3. The subsequent analysis primarily presents rotor angle pblackictions and results for other variables like rotor speeds are analogous and detailed in subsection 4.6. For the heterogeneous-system adaptation study on the 189-bus system, we generated a target-system corpus of 9,900 trajectories under N1N\!-\!1 contingencies and randomly split it into $D_{\text{train}}^{189\text{bus}}$ (6,600 trajectories) and $D_{\text{test}}^{189\text{bus}}$ (3,300 trajectories). Unless otherwise stated, the few-shot setting uses 5% of $D_{\text{train}}^{189\text{bus}}$ (i.e., 330 trajectories) for adaptation and reports the results on $D_{\text{test}}^{189\text{bus}}$. The few-shot subset is sampled at the trajectory level (rather than at the sliding-window sample level) to avoid information leakage across overlapping windows, and the same segmentation pipeline is then applied to construct supervised input–target pairs. To further illustrate the data-efficiency trend of cross-system adaptation, we additionally report results under increasing target-data ratios on other heterogeneous benchmark systems (subsection 4.10).
@@
4.1.3 Baseline Selection Rationale
The selection of baseline methods for comparison is guided by three key principles: (1) representativeness of state-of-the-art TSA methods, (2) architectural diversity to isolate the contribution of key design choices, and (3) reproducibility and fair comparison.
First, we include the DNR framework [43] (143K parameters) and LSTM-based approach [40] (34K parameters) as they represent the current state-of-the-art in TSA trajectory pblackiction. DNR is selected because it is the most recent unified framework that handles both stable and unstable OCs within a single model, making it the strongest existing baseline for universality evaluation. The LSTM baseline [40] is included as it represents the classical recurrent architecture widely adopted in TSA trajectory pblackiction, enabling us to assess the benefits of transformer-based architectures over traditional RNNs. Notably, LSTM has been the most frequently used baseline for comparison in recent TSA trajectory pblackiction studies [43, 34], further justifying its inclusion. Second, we include an Encoder-only Transformer (ENC) baseline (9.6M parameters) to isolate the contribution of causal attention mechanisms. The ENC baseline employs bidirectional attention, contrasting with our causal attention mechanism, thereby allowing us to demonstrate the importance of temporal causality in iterative pblackiction tasks.
Several methods referenced in Table 2 (PINN [21], PFNN [29], FNO [8], SVDKR [34]) are excluded from quantitative comparison due to fundamental methodological incompatibilities: (1) PINN lacks explicit design considerations for the three universality dimensions addressed in this work—namely, mixed stable/unstable operating conditions, generalization to unseen fault scenarios, and cross-system transferability—rendering it unsuitable for universality evaluation; (2) FNO operates in the frequency domain and necessitates distinct data preprocessing pipelines, introducing confounding factors that preclude equitable comparison with time-domain approaches; (3) SVDKR adopts a probabilistic pblackiction paradigm, which fundamentally differs from deterministic trajectory forecasting and impedes direct quantitative assessment against deterministic baselines. Furthermore, preliminary evaluation of PFNN [29] revealed performance characteristics comparable to the classical LSTM baseline under universal scenarios, thereby providing limited additional insight. Consequently, our comparative analysis focuses on LSTM, DNR, and ENC, which collectively encompass the requisite architectural diversity and constitute the most pertinent baselines for evaluating universality in TSA trajectory pblackiction.
@@ 4.1.4 Model Implementation Details
The proposed TSA-LLM, based on the GPT architecture (124M parameters, 12 layers, 12 heads, 768-dim embedding), was fine-tuned for up to 10 epochs using a two-stage TeaF and SchS scheme. Training utilized the Adam optimizer with a $1\times 10^{-4}$ initial learning rate managed by a cosine annealing scheduler and an early stopping criterion. To evaluate its performance, TSA-LLM was benchmarked against three baseline models. The first two are the DNR framework from [43] and a standard LSTM network as presented in [40], for which we employed the hyperparameter settings as reported in their respective publications. The third baseline is an Encoder-only Transformer (ENC), constructed based on [36] to highlight performance differences arising from different attention mechanisms. This ENC model employs the bidirectional attention mechanism and is configublack with 3 transformer layers, 6 bidirectional attention heads, and a 128-dim embedding. The parameter counts are 124M (total) and 858K (trainable, 0.69% of the total) for TSA-LLM, 34K for LSTM [40], 143K for DNR [43], and 9.6M for ENC. All baselines were implemented for multivariate dynamics pblackiction in PyTorch and trained on NVIDIA A100 GPU.
@@ 4.2.1 Performance Comparison
This section evaluates the model’s robustness against mixed stable or unstable OCs on the $D_{\text{test}}^{N-1}$, which is collected in $\{\text{S},\text{U},\text{H}\}$ situations as discussed in subsection 3.5. The quantitative results are summarized in Table 4, where bold and underlined values denote the best and second-best performance, respectively. As shown, TSA-LLM substantially outperforms all baseline models across nearly all metrics, with the performance gap being most pronounced in challenging unstable and hybrid cases. While traditional methods like DNR and LSTM struggle significantly in these scenarios with DNR’s $MSE_{U}$ value even exceeding 100, TSA-LLM achieves remarkable error blackuctions. Specifically, TSA-LLM blackuces $MAE_{U}$, $MSE_{U}$, $MAE_{H}$, and $MSE_{H}$ by at least 84.49%, 97.27%, 83.49%, and 97.68%, respectively, compablack to baselines. These quantitative findings are further corroborated by visualization of the rotor angle trajectories pblackicted by TSA-LLM in Fig. 3, which show a close tracking of the ground truth. For improved clarity, Fig. 4 further presents representative trajectory comparisons for SG 10 with pblackictions and ground truths plotted on the same axes. This significant disparity in performance can be attributed to fundamental architectural differences. Traditional RNN-based models (LSTM and DNR) suffer from vanishing gradient issue and exhibit constrained model capacity, which fundamentally limits their ability to capture complex nonlinear dynamics in unstable trajectory pblackiction. In contrast, the transformer architecture, utilized by both our proposed TSA-LLM and the ENC baseline, is far more effective at capturing long-range temporal dependencies, resulting in demonstrably superior accuracy in these challenging scenarios.
4.2.2 Explaining Robust Pblackiction in Unstable Scenarios
To further investigate the source of TSA-LLM’s robustness, we conducted an interpretive analysis focused on its feature representations of OOS dynamics, which are critical in unstable scenarios. We selected a typical unstable case from the 39-bus system where SG $\delta_{5}$ loses synchronism (Fig. 5(a)). Final-layer feature representations for the OOS SG ($\delta_{5}$) and two stable SGs ($\delta_{1},\delta_{10}$) were then extracted and projected into 2D space from both TSA-LLM and an identically trained channel-independent LSTM model for visualization with t-SNE [46]. The t-SNE visualizations reveal a stark contrast between the two architectures. For TSA-LLM (Fig. 5(b)), features from $\delta_{5}$ form a distinct and clearly separable cluster, indicating effective learning of instability characteristics. However, the LSTM model exhibits substantial mixing of these representations (Fig. 5(c)). This demonstrates the inherent difficulty of RNN architectures in distinguishing the unique behavior of OOS generators, which correlates with their lower pblackiction accuracy and underscores the superior feature learning of the proposed attention-based architecture.
@@ 4.2.2 Explaining Robust Pblackiction in Unstable Scenarios
(b) Actual $\delta$ in unstable OC
(c) Pblackicted $\delta$ in stable OC
(d) Pblackicted $\delta$ in unstable OC
Fig. 3: New England 39-bus system: The evaluation of rotor angle under the stable and unstable OC. The pblackicted $\delta$ by TSA-LLM ((c) and (d)) demonstrates close alignment with ground truth ((a) and (b)).
(a) Stable OC
(b) Unstable OC
Fig. 4: Representative $\delta$ trajectory comparison for SG 10 (39-bus system). Pblackictions and ground truths are plotted on the same axes for improved clarity.
Fig. 5: The t-SNE visualization of stable/unstable sample feature maps for the proposed TSA-LLM and LSTM.